Calculate the mesh currents 11, 12, and iz. Is = 2 A. + 01
Answer #1

from the circuit vo= 2(1,-12) 1 KvL in mesh i -12 + 2 C11-12) +411,-13) = 0 Gil-212 - 413 = 12 ☺ By supermesh Super mesh i3-1Answer question